OpenStudy (anonymous):

Find the x-, y-, and z-intercepts for the equation. 2x - 6y - 3z = 12

hartnn (hartnn):

To get x-intercept put y=0,z=0 in the equation 2x - 6y - 3z = 12 to get y-intercept put x=0,z=0 in the equation 2x - 6y - 3z = 12 to get z-intercept put x=0,y=0 in the equation 2x - 6y - 3z = 12

OpenStudy (anonymous):

oh, okay! that helps so much, thanks :)

hartnn (hartnn):

welcome ^_^